# Reconstruction-Free Anomaly Detection with Diffusion Models

### v2

@@ Abstract
Despite the remarkable success, recent reconstruction-based anomaly detection (AD) methods via diffusion modeling still involve fine-grained noise-strength tuning and computationally expensive multi-step denoising, leading to a fundamental tension between fidelity and efficiency. In this paper, we propose a novel inversion-based AD approach – “detection via noising in latent space” – which circumvents explicit reconstruction. Importantly, we contend that the limitations in prior reconstruction-based methods originate from the prevailing “detection via denoising in RGB space” paradigm. To address this, we model AD under a reconstruction-free formulation, which directly infers the final latent variable corresponding to the input image via DDIM inversion, and then measures the deviation based on the known prior distribution for anomaly scoring. Specifically, in approximating the original probability flow ODE using the Euler method, we only enforce very few inversion steps to noise the clean image to pursue inference efficiency. As the added noise is adaptively derived with the learned diffusion model, the original features for the clean testing image can still be leveraged to yield high detection accuracy. We perform extensive experiments and detailed analysis across three widely used image AD datasets under the unsupervised unified setting to demonstrate the effectiveness of our model, regarding state-of-the-art AD performance, and about $2\times$ inference time speedup without diffusion distillation.
Projecthttps://SkyShunsuke.github.io/InversionADProject/
Codehttps://github.com/SkyShunsuke/InversionAD
Introduction
@@
Image anomaly detection (AD) focuses on identifying images that deviate from normal patterns, with critical applications in industrial defect detection and automated medical diagnosis (J et al. 2020; Bercea et al. 2023). In real-world scenarios, anomalous samples are often scarce and costly to acquire. Consequently, researchers began focusing on unsupervised AD, where the detection model was trained solely on normal images to identify potential anomalies (Roth et al. 2021; Batzner, Heckler, and König 2023).
Due to the robust capabilities for approximating distributions, diffusion models (Ho, Jain, and Abbeel 2020; Sohl-Dickstein et al. 2015) have emerged as powerful tools for AD by precisely modeling normal data distribution (Zhang et al. 2023b; Yao et al. 2024; Lu et al. 2023a; Zhang et al. 2023a; Mousakhan, Brox, and Tayyub 2023; He et al. 2024; Wyatt et al. 2022; Fučka, Zavrtanik, and Skočaj 2025; Teng et al. 2023; Feng et al. 2025; Beizaee et al. 2025). The prevailing paradigm in these methods is to treat anomalies as noise: by adding noise of a certain intensity to an anomalous image and then denoising it with a diffusion model trained solely on normal data, the anomalous regions can be effectively reconstructed as normal regions. The anomaly score is therefore computed by measuring the mean squared error (MSE) between the original input and its reconstructed counterpart, as shown in Figure 1 (a).
We argue that the above “detection via denoising in RGB space” paradigm is less effective given the following fundamental limitations: (1) Noise-strength sensitivity. The detection performance depends heavily on the selection of appropriate noise strengths. Stronger noise corrupts normal regions, increasing false positives, while weaker noise leads to overly well-recovered anomaly regions (Lu et al. 2023a; Yao et al. 2024); (2) Computational expense in multi-step denoising. Satisfactory reconstruction via diffusion models often requires iterative denoising steps, which is significantly time-consuming for real-world applications. As summarized in Table 1, most existing reconstruction-based methods involve high inference latency.
(a) Conventional reconstruction-based AD paradigm
(b) Our inversion-based AD paradigm
Figure 1: Conceptual comparison of conventional and our proposed AD paradigm. and Conventional reconstruction-based paradigm (a) first perturbs an input sample $\mathbf{x}_{0}$ to a latent state $\mathbf{x}_{t}$ at step $t$ , and then denoises $\mathbf{x}_{t}$ back to $\mathbf{x}_{0}$ . The anomaly score is computed as the mean squared error (MSE) between the original input and its reconstructed sample. In contrast, our inversion-based paradigm (b) directly infers the latent state at the final step, $\mathbf{x}_{T}$, by tracing the PF-ODE trajectories. The anomaly score is then determined based on the typicality of $\mathbf{x}_{T}$ within the tractable latent distribution.
Therefore, we introduce the “detection via noising in latent space” paradigm, which bypasses the reconstruction process in prior AD approaches. Our core insight is that, since the learned diffusion model is solely aware of the normal data, we can utilize such learned normal data distribution through latent inversion, rather than reconstruction, as shown in Figure 1 (b). Specifically, for a given image, we directly infer the corresponding latent variable at the final diffusion step by tracing the probability flow ordinary differential equation (PF-ODE (Song et al. 2021)) trajectory. Since PF-ODE is by nature deterministic, this constructs a one-to-one correspondence between the noise prior distribution and the distribution of normal images, ensuring that normal images can be ideally mapped to focus on a high-density region. Conversely, anomalous images will be mapped to low-density regions. Therefore, we can detect anomalies by evaluating the degree of deviation of the inferred latent variable under the prior distribution.
While the above method detours noise strength tuning, it does not straightforwardly lead to a reduction of detection speed. In particular, the high inference latency primarily stems from the high-order ODE solver introduced for accurate reconstruction from inverted latents. To address this, we are inspired by (Teng et al. 2023; Heng, Thiery, and Soh 2024) to measure the norm of the score function as a proxy for typicality, which is robust to the step size selection of the Euler method. This allows us to enforce very few inversion steps to diffuse the clean image to pursue inference efficiency. Moreover, it contributes to a significantly more efficient inference paradigm compared to prior approaches. Importantly, because the learned diffusion model enables noise to be adaptively added during inversion, the key features of the original clean testing image can still be leveraged to harvest high detection accuracy. As our method is designed for the inference stage, it can be applied in a plug-and-play fashion to be incorporated into existing AD models. Experimental evaluations demonstrate that our method substantially improves inference latency, yet still yields high AD performance, across four widely-used AD benchmarks.
Our core contributions include three folds:
  • Proposing a novel paradigm InvAD, detection via noising in latent space, with an inversion-based AD framework that eliminates reconstruction requirements by directly estimating latent variables from inputs and quantifying anomaly scores through deviation measurements under the prior distribution.
  • Demonstrating that the inversion-based approach, which infers latent variables, is simpler, faster, and achieves superior detection performance compared to conventional reconstruction-based methods.
Background
@@
DAD [TPAMI’25] $\{2,400\}$ 9.1 Mask prediction
DDAD [GCPR’24] $\{5,10,25\}$ 0.8 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
ScoreDD [ArXiv’22] 1515 N/A $\nabla_{\mathbf{x}}p_{\theta}(\mathbf{x}_{t})||_{2}$
@@
InvAD (Ours) 33 88.1 $\log p_{\theta}(\mathbf{x}_{T})$
Diffusion Models
Diffusion modeling (Ho, Jain, and Abbeel 2020; Song et al. 2021) is a recently emerged generative modeling framework, which aims to reverse a Markov perturbation process (forward process). In general, each transition of the forward process is defined as:
@@
where monotonically decreasing sequence $\alpha_{1:T}\in(0,1]^{T}$ schedules the noise. The sampling procedure can then be achieved with a reverse process, which defines a Markov process with a Gaussian transition kernel:
@@
where $\mathbf{x}_{t}$ is sampled from $q_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$, and $\mathbf{\gamma}\coloneqq[\gamma_{1},\ldots,\gamma_{T}]^{T}$ denotes the weighting vector. During sampling, one can first sample a pure noise from $p_{\theta}(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$, and then progressively trace Equation 2 for denoising.
Denoising Diffusion Implicit Models
The Gaussian approximation in Equation 2 generally works satisfactorily when the total number of diffusion steps, $T$, is sufficiently large (Ho, Jain, and Abbeel 2020). However, increasing $T$ also magnifies the number of function evaluations (NFE), which slows down the sampling procedure. To address this, Song et al. introduced denoising diffusion implicit models (DDIM), allowing for faster sampling by leveraging a subset of the original diffusion process (Song, Meng, and Ermon 2021). Specifically, they reformulate the reverse process as follows:
@@
Diffusion-Based Anomaly Detection
@@
  • Conditional information injection-based methods incorporate auxiliary information, such as class labels (He et al. 2024), semantic features (Mousakhan, Brox, and Tayyub 2023; He et al. 2024), or unmasked regions (Feng et al. 2025; Beizaee et al. 2025; Xu, Wang, and Du 2024). As pointed out in (Mousakhan, Brox, and Tayyub 2023), conditioning the denosing process with such additional information contributes to reducing reconstruction error in the normal region and increasing it for anomalous regions. However, since this line of methods conditions the diffusion model at training time, they can lead to increased training complexity.
@@
  • Discriminator-based methods introduce an additional discriminator, typically trained to distinguish between normal and anomalous pixels (Zhang et al. 2023b, a). While this line of methods enables efficient anomaly localization, they rely on pseudo-anomalies, biasing AD towards specific types of anomalies.
@@
  • Adaptive timesteps-based methods attempt to adjust or identify perturbation timesteps by comparing the reconstruction errors of denoised results across different time steps (Yao et al. 2024). As the above two types of methods do not incorporate an explicit mechanism for adjusting the perturbation time step, it is typically treated as a hyperparameter or addressed via ensembling over multiple time steps (Lu et al. 2023a). Details can be found in the ”timestep tuning” column of Table 1.
@@
In contrast to the state-of-the-art non-diffusion method (Batzner, Heckler, and König 2023) that achieves over 200 FPS, current diffusion-based methods mostly reach 1 FPS on average. While a recent diffusion-based approach (Feng et al. 2025) tackles this issue by introducing score distillation to constrain the diffusion iteration to only one step, the training for distillation can be complicated due to the instability of the adversarial learning. Table 1 summarizes the properties of current diffusion-based AD methods. Overall, a simple, efficient, yet effective paradigm for diffusion-based AD remains unexplored to date.
Methodology
@@
To address the limitations of current diffusion-based methods, we introduce a novel diffusion-based AD framework with latent inversion, where the sampling trajectory is reformulated as $\mathbf{x}_{0}\rightarrow\mathbf{x}_{T}$. In contrast to conventional reconstruction-based approaches $\bigl{(}\mathbf{x}_{0}\rightarrow\mathbf{x}_{t}\rightarrow\hat{\mathbf{x}}_{0}\bigr{)}$, our method infers the final latent state $\mathbf{x}_{T}$ directly from $\mathbf{x}_{0}$ by following the learned PF-ODE trajectory. Figure 1(b) illustrates our reconstruction-free AD paradigm, which leverages the inversion process to achieve efficient yet accurate AD.
Inversion-Based Anomaly Detection
@@
Given an image $\mathbf{x}$ and a score function $\epsilon_{\theta}^{(t)}$ trained using Equation 3, our goal is to infer the latent representation $\mathbf{x}_{T}$ corresponding to $\mathbf{x}$ by tracing the PF-ODE trajectory. In this work, we employ the PF-ODE derived from the DDIM framework (Song, Meng, and Ermon 2021), as shown in Equation 5. By applying an Euler approximation to Equation 5 and integrating forward, we obtain the following discrete update equation for latent inversion:
@@
Intuitively, Equation 7 leads to a considerable speed-up for the inference stage when $T\gg S$. Nevertheless, it also inevitably sacrifices approximation precision as the step interval for the Euler integration is extended. Recall here that our goal is to resolve AD rather than pursue accurate pixel-wise reconstruction. As such, we argue that high AD performance does not necessarily require reconstruction precision. It should be noticed that, even with limited inversion steps, the inversion mechanism in projecting an anomaly pixel can nonetheless yield sufficiently low typicality. In this regard, despite the poorly reconstructed samples, an anomaly pixel can still be expectedly mapped to a low-density region within the noise prior to achieve high detection performance. This insight suggests that our inversion-based AD method is free from the accuracy-efficiency trade-off in conventional reconstruction-based methods. We next need to find an ideal space that best suits our diffusion-based AD modeling.
@@
Diffusion Modeling in Feature Space
Current image AD methods (Batzner, Heckler, and König 2023; Roth et al. 2021; Um et al. 2023; Mousakhan, Brox, and Tayyub 2023) heavily rely on pre-trained networks, referred to as backbone models, that are trained on large-scale datasets such as ImageNet (Deng et al. 2009). In diffusion-based AD, these backbones are used to compare the original test image with its reconstructed counterpart in feature space (Mousakhan, Brox, and Tayyub 2023; Lu et al. 2023a; He et al. 2024). Because backbone features exhibit invariance to low-level variations (e.g., noise or lighting conditions), they can significantly enhance AD performance.
Table 2: Quantitative results on different AD datasets under the multi-class setting. The best and the second-best results are highlighted in bold and underlined, respectively. For each row, the mAD averages the evaluation over all image- and pixel-level metrics (i.e., the 3rd to the 10th column). FPS measures the image-level detection efficiency.

### v3

# InvAD: Inversion-based Reconstruction-Free Anomaly Detection with Diffusion Models

@@ Abstract
Despite the remarkable success, recent reconstruction-based anomaly detection (AD) methods via diffusion modeling still involve fine-grained noise-strength tuning and computationally expensive multi-step denoising, leading to a fundamental tension between fidelity and efficiency. In this paper, we propose InvAD, a novel Inversion-based Anomally Detection approach – “detection via noising in latent space” – which circumvents explicit reconstruction. Importantly, we contend that the limitations in prior reconstruction-based methods originate from the prevailing “detection via denoising in RGB space” paradigm. To address this, we model AD under a reconstruction-free formulation, which directly infers the final latent variable corresponding to the input image via DDIM inversion, and then measures the deviation based on the known prior distribution for anomaly scoring. Specifically, in approximating the original probability flow ODE using the Euler method, we only enforce very few inversion steps to noise the clean image to pursue inference efficiency. As the added noise is adaptively derived with the learned diffusion model, the original features for the clean testing image can still be leveraged to yield high detection accuracy. We perform extensive experiments and detailed analysis across four widely used industrial and medical AD benchmarks under the unsupervised unified setting to demonstrate the effectiveness of our model, achieving state-of-the-art AD performance, and about $2\times$ inference time speedup without diffusion distillation. Code is available at https://github.com/SkyShunsuke/InversionAD.
1 Introduction
@@
Image anomaly detection (AD) focuses on identifying images that deviate from normal patterns, with critical applications in industrial defect detection and automated medical diagnosis [Yang2020, bercea2023generalizing]. In real-world scenarios, anomalous samples are often scarce and costly to acquire. Consequently, researchers began focusing on unsupervised AD, where the detection model was trained solely on normal images to identify potential anomalies [PatchCore, EfficientAD].
Figure 1: Accuracy v.s. Speed relationship of diffusion-based AD methods [DiAD, TransFusion, omiad, decodiff, mad] on MVTecAD. Our proposed InvAD achieves state-of-the-art AD performance with a substantially speedup.
@@
MDM [ICML’25] 4040 1.9 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
@@
InvAD (Ours) 3 88.1 $\log p_{\theta}(\mathbf{x}_{T})$
Due to the robust capabilities for approximating distributions, diffusion models [DDPM, DiffusionModel, dhariwal2021diffusion] have emerged as powerful tools for AD by precisely modeling normal data distribution [DiffAD, GLAD, RemoveAsNoise, DiffusionAD, DDAD, DiAD, AnoDDPM, TransFusion, ScoreDD, omiad, decodiff, dte]. The prevailing paradigm in these methods is to treat anomalies as noise: by adding noise of a certain intensity to an anomalous image and then denoising it with a diffusion model trained solely on normal data, the anomalous regions can be effectively reconstructed as normal regions. The anomaly score is therefore measured with the mean squared error (MSE) between the original input and its reconstructed counterpart, as shown in Figure 2 (a).
We argue that the above “detection via denoising in RGB space” paradigm is less effective given the following fundamental limitations: (1) Noise-strength sensitivity. The detection performance depends heavily on the selection of appropriate noise strengths. Stronger noise unnecessarry corrupts normal regions, increasing false positives, while weaker noise leads to overly well-recovered anomaly regions [RemoveAsNoise, GLAD]; (2) Computational expense in multi-step denoising. Satisfactory reconstruction via diffusion models often requires iterative denoising, which is time-consuming for real-world applications. As listed in Table 1, most reconstruction-based methods involve high inference latency.
Therefore, we introduce the “detection via noising in latent space” paradigm, which bypasses the reconstruction process in prior AD approaches. Our core insight is that, since the learned diffusion model is solely aware of the normal data, we can utilize such learned normal data distribution through latent inversion, rather than reconstruction, as shown in Figure 2 (b). Specifically, for a given image, we directly infer the corresponding latent variable at the final diffusion step by tracing the probability flow ordinary differential equation (PF-ODE [SBGM]) trajectory. Since PF-ODE is by nature deterministic, this constructs a one-to-one correspondence between the noise prior distribution and the distribution of normal images, ensuring that normal images can be ideally mapped to focus on a high-density region. Conversely, anomalous images will be mapped to low-density regions. Therefore, we can detect anomalies by evaluating the degree of deviation of the inferred latent variable under the prior distribution.
(a) Conventional reconstruction-based AD paradigm
(b) Our inversion-based AD paradigm
Figure 2: Conceptual comparison of conventional and our proposed AD paradigm. Conventional reconstruction-based paradigm (a) first perturbs an input sample $\mathbf{x}_{0}$ to a latent state $\mathbf{x}_{t}$ at step $t$ , and then denoises $\mathbf{x}_{t}$ back to $\mathbf{x}_{0}$ . The anomaly score is computed as the mean squared error (MSE) between the original input and its reconstructed sample. In contrast, our inversion-based paradigm (b) directly infers the latent state at the final step, $\mathbf{x}_{T}$, by tracing the PF-ODE trajectories. The anomaly score is then determined based on the typicality of $\mathbf{x}_{T}$ within the tractable latent distribution.
While the above method detours noise strength tuning, it does not straightforwardly lead to a reduction of detection speed. In particular, the high inference latency primarily stems from the high-order ODE solver introduced for accurate reconstruction from inverted latents. To address this, we are inspired by [ScoreDD, diffpath] to measure the norm of the score function as a proxy for typicality, which is robust to the step size selection of the Euler method. This allows us to enforce very few inversion steps to diffuse the clean image to pursue inference efficiency. Moreover, it contributes to a significantly more efficient inference paradigm compared to prior approaches. Importantly, because the learned diffusion model enables noise to be adaptively added during inversion, the key features of the original clean testing image can still be leveraged to harvest high detection accuracy. As plotted in Figure 1, our InvAD achieves state-of-the-art detection performance with the inference efficiency surpassing prior diffusion-based AD models by a large margin. Also, as our method is designed for the inference stage, it can be applied in a plug-and-play fashion to be incorporated into existing AD models. Experimental evaluations demonstrate that our method substantially improves inference latency, yet still yields high AD performance, across four widely-used AD benchmarks.
Our core contributions include three folds:
  • Proposing a novel paradigm, detection via noising in latent space, with an inversion-based AD framework InvAD that eliminates reconstruction requirements by directly estimating latent variables from inputs and quantifying anomaly scores through deviation measurements under the prior distribution.
  • Demonstrating that our inversion-based approach, which infers latent variables, is simpler, faster, and achieves state-of-the-art detection performance compared to conventional reconstruction-based methods.
2 Background
2.1 Diffusion Models
Diffusion modeling [DDPM, SBGM] is a recently emerged generative modeling framework, which aims to reverse a Markov perturbation process (forward process). In general, each transition of the forward process is defined as:
@@ 2.1 Diffusion Models
where the monotonically decreasing sequence $\alpha_{1:T}\in(0,1]^{T}$ schedules the noise intensity. The sampling procedure can then be achieved with a reverse process, which defines a Markov process with a Gaussian transition kernel:
@@
where $\mathbf{x}_{t}$ is efficiently sampled from $q_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$, and $\mathbf{\gamma}\coloneqq[\gamma_{1},\ldots,\gamma_{T}]^{T}$ denotes the weighting vector. During sampling, one can first sample a pure noise from $p_{\theta}(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$, and then progressively trace Equation 2 for denoising.
2.2 Denoising Diffusion Implicit Models
The Gaussian approximation in Equation 2 generally works satisfactorily when the total number of diffusion steps, $T$, is sufficiently large [DDPM]. However, increasing $T$ also magnifies the number of function evaluations (NFE), which slows down the sampling procedure. To address this, Song et al. introduced denoising diffusion implicit models (DDIM), allowing for faster sampling by leveraging a subset of the original diffusion process [DDIM]. Specifically, they reformulate the reverse process as follows:
@@ 2.2 Denoising Diffusion Implicit Models
2.3 Diffusion-Based Anomaly Detection
@@
  • Conditional information injection-based methods incorporate auxiliary information, such as class labels [DiAD], semantic features [DDAD, DiAD], or unmasked regions [omiad, maediff, mad]. As pointed out in [DDAD], conditioning the denosing process with such additional information contributes to reducing reconstruction error in the normal region and increasing it for anomalous regions. However, since this line of methods conditions the diffusion model at training time, they can lead to increased training complexity.
@@
  • Discriminator-based methods introduce an additional discriminator, typically trained to distinguish between normal and anomalous pixels [DiffAD, DiffusionAD]. While this line of methods enables efficiency, they rely on pseudo-anomalies, biasing AD towards specific anomaly types.
@@
  • Adaptive timesteps-based methods attempt to adjust or identify perturbation timesteps by comparing the reconstruction errors of denoised results across different time steps [GLAD]. As the above two types of methods do not incorporate an explicit mechanism for adjusting the perturbation time step, it is typically treated as a hyperparameter or addressed via ensembling over multiple time steps [RemoveAsNoise]. See the 5th column in Table 1 for more details.
@@
In contrast to the state-of-the-art non-diffusion method [EfficientAD] that achieves over 200 FPS, current diffusion-based methods mostly reach 1 FPS on average. While a recent diffusion-based approach [omiad] tackles this by introducing score distillation to constrain the diffusion iteration to only one step, the training for distillation can be complicated due to the instability of the adversarial learning. Table 1 summarizes the properties of current diffusion-based AD methods. Overall, a simple, efficient, yet effective paradigm for diffusion-based AD remains unexplored to date.
3 Methodology
@@
To address the limitations of current diffusion-based methods, we introduce a novel diffusion-based AD framework with latent inversion, where the sampling trajectory is reformulated as $\mathbf{x}_{0}\rightarrow\mathbf{x}_{T}$. In contrast to conventional reconstruction-based approaches $\bigl(\mathbf{x}_{0}\rightarrow\mathbf{x}_{t}\rightarrow\hat{\mathbf{x}}_{0}\bigr)$, our method infers the final latent state $\mathbf{x}_{T}$ directly from $\mathbf{x}_{0}$ by following the learned PF-ODE trajectory. Figure 2(b) illustrates our reconstruction-free AD paradigm, which leverages latent inversion to achieve efficient yet accurate AD.
3.1 Inversion-Based Anomaly Detection
@@
Given an image $\mathbf{x}$ and a score function $\epsilon_{\theta}^{(t)}$ trained using Equation 3, our goal is to infer the latent representation $\mathbf{x}_{T}$ corresponding to $\mathbf{x}$ by tracing the PF-ODE trajectory. In this work, we employ the PF-ODE derived from the DDIM framework [DDIM], as shown in Equation 5. By applying an Euler approximation to Equation 5 and integrating forward, we obtain the following discrete update equation for latent inversion:
@@
Intuitively, Equation 7 leads to a considerable speed-up for the inference stage when $T\gg S$. Nevertheless, it also inevitably sacrifices approximation precision as the step interval for the Euler integration is extended. Recall here that our goal is to resolve AD rather than pursue accurate pixel-wise reconstruction. As such, we argue that high AD performance does not necessarily require reconstruction precision. It should be noticed that, even with limited inversion steps, the inversion mechanism in projecting an anomaly pixel can nonetheless yield sufficiently low typicality. In this regard, despite the poorly reconstructed samples, an anomaly pixel can still be expectedly mapped to a low-density region within the noise prior to achieve high detection performance. This insight suggests that our inversion-based AD method is free from the accuracy-efficiency trade-off in conventional reconstruction-based methods. We next need to find an ideal space that best suits our diffusion-based AD modeling. See Supp. Mat. for more detail deriviation of an inversion fomula and analysis on inversion.
3.2 Diffusion Modeling in Feature Space
Current image AD methods [EfficientAD, PatchCore, FastFlow, DDAD] heavily rely on pre-trained networks, referred to as backbone models, that are trained on large-scale datasets such as ImageNet [ImageNet]. In diffusion-based AD, these backbones are used to compare the original test image with its reconstructed counterpart in feature space [DDAD, RemoveAsNoise, DiAD]. Because backbone features exhibit invariance to low-level variations (e.g., noise or lighting conditions), they can significantly enhance AD performance.
@@ 3.2 Diffusion Modeling in Feature Space
3.3 Anomaly Score Calculation
Once we obtain the latent representation $\mathbf{z}_{T}\in\mathbb{R}^{C\times h\times w}$ for features of a given test sample $\mathbf{x}$, the most straightforward way to compute the anomaly score is via the log-likelihood under the latent Gaussian distribution. However, it often triggers the known reverse-scoring problem and fails to capture small anomalies (See Supp. Mat.). Given that anomalies are inherently sparse in images, we additionally propose a norm-based image-level anomaly scoring approach. Specifically, we compute the Euclidean norm over all channels $C$ for each pixel $(i,j)$ in $\mathbf{z}_{T}$, which is written as
Different from [DiAD], for the image-level anomaly score, we compute the absolute difference between the maximum and minimum values of $\mathbf{z}_{T}^{\text{normed}}$. This is motivated by the observation that, even if an image is anomalous, anomalies for images are often locally and sparsely distributed within small regions, and the majority of image regions remain normal. In essence, our scheme for image-level anomaly scoring examines the difference between the highest- and the lowest-norm pixels to mitigate the influence of potential outliers, thereby alleviating the reverse-scoring issue. The pixel-level anomaly score map $M_{\text{px}}$ is simply produced by interpolating $\mathbf{z}_{T}^{\text{normed}}\in\mathbb{R}^{h\times w}$ to match the desired output resolution $(H,W)$.
4 Experiments
4.1 Datasets
We use three widely-used unsupervised AD benchmarks—MVTecAD [MVTecAD], VisA [ViSA], MPDD [MPDD]—all designed for measuring real-world AD performance on RGB industrial images. We also experiment on the BMAD [bmad] benchmark (with six reorganized datasets) to study the effectiveness of InvAD on the medical AD domain. Since each benchmark provides both image- and pixel-level annotations, we can evaluate performance in terms of overall AD (image-level) as well as precise localization (pixel-level). See Supp. Mat. for a detailed description of these datasets.
Table 2: Quantitative results on different AD datasets under the multi-class setting. The best and the second-best results are highlighted in bold and underlined, respectively. For each row, the mAD averages the evaluation over all image- and pixel-level metrics (i.e., the 3rd to the 9th column). FPS measures the detection efficiency.